OpenStudy (anonymous):

the reflecting dish of a parabolic microphone has a cross-section in the shape of a parabola.the microphone itself is placed on the focus of the parabola.if the parabola is 40 inches wide and 20 inches deep,how far from the vertex should the microphone be placed?

OpenStudy (cwrw238):

this is a parabola y^2 = 4ax with focus at (a,0) so plugging in corresponding values of x and y 20^2 = 4a * 20 = 80a giving a = 400/80 = 5 microphone must be placed 5 inches form the vertex

OpenStudy (anonymous):

thank you so much<3

OpenStudy (cwrw238):

yw